# Improving Spiking Neural Network Accuracy With Color Model Information Encoded Bit Planes

### v1

@@ Abstract
Spiking neural networks (SNNs) have emerged as a promising paradigm in computational neuroscience and artificial intelligence, offering advantages such as low energy consumption and small memory footprint. However, their practical adoption is constrained by several challenges, prominently among them being performance optimization. In this study, we present a novel approach to enhance the performance of SNNs through a new encoding method that exploits bit planes derived from various color models of input image data for spike encoding. Our proposed technique is designed to improve the computational accuracy of SNNs compared to conventional methods without increasing model size. Through extensive experimental validation, we demonstrate the effectiveness of our encoding strategy in achieving performance gain across multiple computer vision tasks. To the best of our knowledge, this is the first research endeavor applying color spaces within the context of SNNs. By leveraging the unique characteristics of color spaces, we hope to unlock new potentials in SNNs performance, potentially paving the way for more efficient and effective SNNs models in future researches and applications.
@@
Spiking neural network, spike encoding, image classification
@@ I Introduction
The field of artificial intelligence (AI) has experienced unprecedented advancements in recent decades[1], largely driven by the development and optimization of deep learning algorithms. Despite these advancements, traditional artificial neural networks (ANNs) often fall short in modeling the intricate and dynamic nature of biological neural systems[2]. To bridge this gap, spiking neural networks (SNNs) have emerged as a promising alternative, leveraging the temporal dimension of neuronal activity to more closely emulate the behavior of biological brains.
@@
Prior research in ANNs has extensively explored the application of image color spaces to enhance performance across various tasks, demonstrating significant improvements in areas such as image recognition and processing[10, 11]. Motivated by these advancements, we aim to investigate whether these insights on color spaces can similarly improve the performance of SNNs. This led us to formulate our research questions:
  • Does encoding color spaces encapsulate the necessary information to enhance the performance of SNNs?
  • How can we encode color spaces information into SNNs?
To answer these hypotheses, we observed that the bit planes of images contain pulse-like information, which has been utilized previously on ANNs [12] and can be apply to encode color spaces information for SNNs to possibly improve accuracy.
@@ II Spiking Neural Network
In this section we will cover some literature survey related to our research, particularly about SNNs in general and some common knowledge on SNN architecture utilized in our experiments.
@@ II-B The Integrate-and-Fire model
While previous efforts have successfully converted ANN neurons with ReLU activations to IF neurons, challenges still remain in the conversion process. One of the key issues is the difficulty in precisely determining the appropriate thresholds in these conversion approaches. Despite the introduction of both data-driven and model-driven threshold determination methods[21], the converted SNN often experience unstable accuracy compared to their re-trained ANN counterparts. Additionally, IF neurons lack time-dependent memory. Once a neuron receives sub-threshold potentials, the membrane potential persists without decay until the neuron fires. This behavior is said to be inconsistent with biological neurons, which exhibit dynamic temporal properties [22].
@@ II-C Deep Residual Learning in Spiking Neural Networks
To address these issues, SEW ResNet[25] was introduced as a replacement design for residual blocks while utilizing IF neurons. The SEW block capitalizes on the binary nature of spikes by applying element-wise functions $g$ to the output of the residual mapping $A_{l}[t]$ and the input spike train $S_{l}[t]$[25]. This approach facilitates identity mapping, mitigates the vanishing/exploding gradient problem and achieves higher accuracy with fewer time steps compared to existing SNN models. With these advantages in mind, we have decide to choose (but not limited to) SEW Resnet[25] architecture as our main experimenting architecture.
III Spike Coding in SNNs
In the context of SNN there are several encoding methods, each with unique advantages and disadvantages.
  • Rate coding [26, 27] is the predominant coding technique utilized for spikes in SNNs. In rate coding, the number of spikes within a specific time interval is counted, and the firing rate is interpreted as the signal intensity. This method inherently includes a high level of redundancy and exhibits tolerance to a certain degree of noise.
  • Temporal coding[3] is the form of coding where spike patterns are distributed over time. Population-based coding methods [28] encode information based on the number of spikes within a given time frame. In contrast, inter-spike interval methods [29] use the time between two spikes as their coding scheme. Temporal coding allows for arbitrarily sized data patterns, though longer patterns require extended time periods.
We primarily focus on rate coding, as it is one of the most widely used encoding schemes in conventional SNNs for representing signals with spike trains. In the rate coding scheme, the firing frequency of a neuron represents the signal intensity. The firing rate $r_{i}$ of the i-th neuron provided by [30] can be expressed as follows:
$r_{i}=\frac{N_{s}}{T}=\frac{\sum_{t=1}^{T}s_{i}(t)}{T}$ (2)
When $N_{s}$ spikes are fired over T time steps, each time step can either produce a spike or none. The firing rate is 1 if a neuron fires at every time step, and 0 if it does not fire at all. In image classification, for instance, if a pixel intensity ranges [0.0, 1.0], it can be represented by a spike train with a corresponding firing rate [0.0, 1.0] as input to the SNN.
In this research, we will be using the Poisson-distributed rate coding [30, 31] scheme where spike $s_{i}$ for i-th pixel with intensity $P_{i}$ can be interpreted as:
@@ III Spike Coding in SNNs
$s_{i}=\begin{cases}1&\text{if $random(0.0,1.0)\geq P_{i}$}\\ 0&\text{otherwise}\end{cases}$ (3)
IV Color spaces and their application in deep learning
@@
This section will provide an overview of relevant concepts for our research, focusing specifically on the color models explored and their applications in deep learning algorithms.
IV-A Color spaces
Color spaces are fundamental in computer graphics as they define how colors are represented, manipulated, and displayed within digital images. In computer graphics applications, understanding and effectively managing color spaces is crucial for tasks such as rendering realistic scenes, creating digital artwork, and designing user interfaces. In this subsection, we will discuss the color spaces under investigation in our experiments, divided by certain criteria from [32, 33], including: device-oriented models, user-oriented models and device-independent models.
IV-A1 Device-oriented models
These models are influenced by the signals of the devices and the resulting colors are affected by the tools used for display. Widely utilized in various applications that require color consistency with hardware tools, examples include any hardware devices used for human visual perception, such as televisions and video systems . While there are different models for separate use-cases and devices, for now we will only concern with RGB, CMY(K) and YCbCr.
IV-A2 User-oriented models
User-oriented color models are considered as a pathway between the observer and the device handling the color information. These models enable users to describe and approximate their color perceptions from the presented colors. For our experiments, we mostly investigate cylindrically described models, including HSL and HSV. According to [32], these models also inherit other special properties such as devices dependence and cylindrical-coordinate representation.
IV-A3 Device-independent models
Device-independent color models are essential for defining color signals without reliance on the specific characteristics of any particular device or application. These models play a crucial role in applications that involve color comparisons and the transmission of visual information across networks that connect various hardware platforms. We will mostly research color spaces defined by the International Commission on Illumination (CIE), including CIE XYZ and CIE LAB.
IV-B Usage of color models in deep learning
Color models are becoming increasingly important in deep learning algorithms, offering valuable information that enhances model performance in various tasks. For example, research on traffic light recognition[34] tested a range of color models across different deep learning architectures, with results indicating that the RGB model remains favored despite some of its limitations. This suggests that, while RGB is not perfect, it still outperforms other models in certain applications.
In another study focused on applying convolutional neural networks (CNNs) for skin detection against adversarial attacks of generative models[35], RGB and YCbCr color models demonstrated superior performance, particularly in handling varying lighting conditions. This highlights the adaptability of these color models in scenarios where lighting inconsistencies can degrade model accuracy.
Moreover, the development of ColorNet[10] illustrates the potential of leveraging multiple color spaces to capture a broader set of features. By combining information from these diverse color spaces, ColorNet outperforms traditional models that rely on a single color space. However, this approach introduces computational overhead due to the inclusion of smaller submodels, which can be a challenge despite the model’s increased efficiency in some cases.
Despite extensive researches on the use of color models in ANNs, their application in SNNs remains underexplored and requires further investigation.
V Bit planes application in deep learning
@@
Bit planes have shown extensive utility in deep learning applications, offering new ways to improve model efficiency and performance. In one study, researchers used bit planes to encode the input layer in Binary Neural Networks (BNNs)[12]. They argued that bit planes retain crucial spatial information, enabling a reduction in the number of multiply-accumulate operations (MACs) without compromising the model’s accuracy or robustness. This demonstrates the potential of bit planes in streamlining computations while maintaining competitive performance.
@@
Another study explored the use of bit plane slicing on RGB images for adversarial robustness[36]. By selecting the most significant bit planes, the researchers developed a classification model that was better equipped to handle adversarial attacks. This approach capitalizes on the spatial information encoded in different bit planes, which helps the model resist manipulations that would otherwise degrade its performance.
@@
Similar to color models, although bit planes have been widely studied and applied in ANNs, their utilization in SNNs is still relatively unexplored and demands further research.
VI Experiments setup
In this section, we will discuss in details about the implementation of our encoding algorithm and setup for experimenting with this method.
VI-A Preliminary testing
As theoretically mentioned in [38] and practically measured in [39, 40], a Poisson-distributed time steps appear to positively correlate with training accuracy in SNNs and to achieve a correct spike distribution, the input must be sampled to a certain frequency. However in the case of a high time steps input, this approach sacrifices the energy efficiency of SNNs, which is one of their primary advantages. Despite this trade-off, the improvement in training accuracy suggests that the benefits in performance might outweigh the costs in specific applications where accuracy is paramount.
@@ VI-A Preliminary testing
We previously hypothesize that the information contained in the bit planes of color models can be effectively used to encode inputs for SNN training. Bit planes decompose an image into binary levels, each representing a specific bit of the pixel values. This decomposition can potentially preserve critical information from the original image while providing a format that might be more suitable for the spike-based processing in SNNs. The granularity of information available in bit planes offers a unique approach to encoding visual data, which could enhance the learning capabilities of SNNs by providing a richer, multi-level representation of the input.
To test our hypotheses, we conducted preliminary experiments using the CIFAR-10 dataset, employing the bit plane encoding method. We compared this approach with traditional rate coding and a combination of both methods (illustrated in details in Figure 2 and Section VI-B) with a rate coding step size $T=10$ (illustrated in Figure 1 and Table I).
Figure 1: Compare training progression on CIFAR-10 [41] between using traditional rate coding (denoted blue), encoding using only bit planes information (denoted orange) and combining bit planes information with rate coding (denoted green).
Encoding method Validation Accuracy (%)
Poisson rate coding with time step T=10 70.69
RGB bit plane encoding 73.49
Poisson rate coding with time step T=10 and combined RGB bit plane encoding 37.28
TABLE I: Validation accuracy comparison on CIFAR-10 [41] between traditional rate coding, encoding using only bit planes information and combining bit planes information with rate coding.
Visible from our results, feeding bit plane information of the encoded RGB color model to the SNNs allowed the model to converge to a certain level of accuracy. Still, the accuracy of purely bit planes encoded inputs still fall short compare to rate coding. On the other hand, a combination of rate coding and bit planes encoding experience an increase in performance, this outcome demonstrates that bit planes in color models contain valuable information that SNNs can learn from to improve their classification ability and encourage us to perform testing on other datasets with different color models (as mentioned in Section IV-A).
VI-B Overall encoding scheme with color spaces
@@
Figure 2: Our proposed encoding method (lower) compare with traditional rate coding method (upper).
@@
As we can see in Figure 2, for our research the encoding scheme adopted integrates a dual approach featuring both traditional rate coding and bit plane encoding methodologies. Initially, the process commences with the conventional rate coding procedure applied to the image tensor. Subsequently, depending on the chosen color model (for example RGB), the image tensor may undergo conversion to the specified model or proceed directly to encoding if already in RGB format. Following this, the bit plane encoding algorithm (will be discussed in details in Section VI-C and Algorithm 1) is employed to transform the color-modeled image into binary planes. The final encoded tensors are then generated through tensor concatenation across the first dimension, merging the outputs from rate coding and bit plane encoding. These encoded tensors are thereby prepared for utilization in inference tasks and/or training processes, facilitating enhanced performance in our computational models.
VI-C Bit plane encoding algorithm
@@
To achieve our research objectives, we have developed Algorithm 1 designed to encode an image tensor batch $X$ into $n_{bit}$ bit planes. To determine $n_{bit}$, it is essential to know the maximum possible value $X_{max}$ of the tensor batch. Since different color models have varying $X_{max}$, this value must be predefined according to the selected color model (for example, if RGB is chosen, then $X_{max}=255$). Once $X_{max}$ is defined, $n_{bit}$ can be calculate as $\left\lceil\log_{2}(X_{max})\right\rceil$.
@@
To calculate the bit planes of an unnormalized integer input tensor batch $X$, we perform element-wise division of $X$ by 2. The remainder from this division constitutes a bit plane of $X$, while the quotient is saved for the extraction of subsequent bit planes. This process is repeated until we obtain $n_{bit}$ bit planes. The resulting bit planes are then concatenated to form a binary tensor batch with the shape $(n_{bit},B,C,H,W)$. This format closely resembles that of rate encoding and ensures determinism in the encoding results without having to sample more than $n_{bit}$ times.
Algorithm 1 Bit plane encoding algorithm
$X_{shape}\in(B,C,H,W)$triangleright Constraining input format of multiple images. If input is a single image we should have $B=1$.
$X_{max}$ is defined
function BiplaneEncode($X,X_{max}$)
@@
     $n_{bit}\leftarrow\left\lceil\log_{2}(X_{max})\right\rceil$triangleright Get number of bit required to encode image from highest possible value $X$.
     $L\leftarrow[\text{ }]$triangleright Empty list initialization.
@@
         $L.insert(X\mod{2})$triangleright Element-wise modulo of tensor, then store the bit plane.
         $X\leftarrow\left\lfloor X/2\right\rfloor$triangleright Element-wise division of tensor.
@@
     return $concat(L)$triangleright Return a concatenated tensor batches of new dimension $(n_{bit},B,C,H,W)$
@@
VI-D Experiment setup
To investigate performance gains during training process, the experimented model should be a medium-sized SNN with potential for improvement in computer vision tasks. After careful consideration, we had decided to select SEW-ResNet18 with ADD element-wise function [25] which is a novel medium size SNN tested on various computer vision tasks and suitable to our specifications. The model and experiment process was implemented using PyTorch [42] and SpikingJelly[43], incorporating color conversion algorithms based on implementations from Kornia [44]. During the rate coding steps, all training processes utilized step size $T=10$. For optimization, we employed the Adam optimizer [45] with a learning rate of $lr=1e-3$ and $\beta=(0.9,0.999)$ for all tested datasets. Cross-Entropy Loss with Softmax function and mean reduction $L(x,y)$[42] across $N$ mini-batch dimension was also widely employed as the loss function:

### v2

# Improvement of Spiking Neural Network with Bit Planes and Color Models

@@ Abstract
Spiking neural network (SNN) has emerged as a promising paradigm in computational neuroscience and artificial intelligence, offering advantages such as low energy consumption and small memory footprint. However, their practical adoption is constrained by several challenges, prominently among them being performance optimization. In this study, we present a novel approach to enhance the performance of SNN for images through a new coding method that exploits bit plane representation. Our proposed technique is designed to improve the accuracy of SNN without increasing model size. Also, we investigate the impacts of color models of the proposed coding process. Through extensive experimental validation, we demonstrate the effectiveness of our coding strategy in achieving performance gain across multiple datasets. To the best of our knowledge, this is the first research that considers bit planes and color models in the context of SNN. By leveraging the unique characteristics of bit planes, we hope to unlock new potentials in SNNs performance, potentially paving the way for more efficient and effective SNNs models in future researches and applications.
@@
Spiking neural network, spike coding, image classification
@@ I Introduction
The field of artificial intelligence (AI) has experienced unprecedented advancements in recent decades[1], largely driven by the developments of deep learning algorithms. Despite these advancements, traditional artificial neural networks (ANNs) often fall short in modeling the intricate and dynamic nature of biological neural systems[2]. To bridge this gap, spiking neural networks (SNNs) have emerged as a promising alternative, leveraging the temporal dimension of neuronal activity to more closely emulate the behavior of biological brains.
@@
Prior research in ANN has explored the application of bit planes (e.g. [10]) and color models [11, 12] to enhance performance across various tasks, demonstrating significant improvements in areas such as image recognition and processing. Motivated by these advancements, we aim to investigate whether bit planes and color models can similarly improve the performance of SNNs. This led us to formulate our research questions:
  • Does bit plane representation enhance the performance of SNN?
  • What are impacts of different color models on SNN?
Especially, we observe that bit planes of an image contain pulse-like information, and can be applied to encode color information for SNNs to possibly improve accuracy.// After extensive experimentation, we concluded that bit planes contain valuable information that can be utilized in the rate coding process as a supplementary source to enhance the accuracy of direct surrogate backpropagation process [13] in SNN architectures. Furthermore, bit plane information proved to be reusable across multiple color models in most of the cases examined in our study.
A preliminary version of this paper is published in [14], where the benefit of including bit plane coding is proved by three grayscale datasets. In this extended version, the following new points have been added. First, the proposed method was updated to support color images. Second, eight more datasets are included. Third, impacts of color models are investigated.
@@ II Spiking Neural Network
In this section we present an overview of SNN knowledge related to our research.
@@ II-B The Integrate-and-Fire model
While previous efforts have successfully converted ANN neurons with ReLU activations to IF neurons, challenges still remain in the conversion process. One of the key issues is the difficulty in precisely determining the appropriate thresholds in these conversion approaches. Despite the introduction of both data-driven and model-driven threshold determination methods[23], the converted SNN often experience unstable accuracy compared to their re-trained ANN counterparts. Additionally, IF neurons lack time-dependent memory. Once a neuron receives sub-threshold potentials, the membrane potential persists without decay until the neuron fires. This behavior is said to be inconsistent with biological neurons, which exhibit dynamic temporal properties [24]. Still, within certain SNN architectures, IF neurons have demonstrated superior performance compared to the supposedly better LIF neurons[24], with the SEW-ResNet[25] architecture being one such example.
Figure 1: Illustrated bit planes of a sample image from CIFAR10[26] dataset.
@@ II-C Deep Residual Learning in Spiking Neural Networks
To address these issues, SEW ResNet[25] was introduced as a replacement design for residual blocks while utilizing IF neurons. The SEW block capitalizes on the binary nature of spikes by applying element-wise functions $g$ to the output of the residual mapping $A_{l}[t]$ and the input spike train $S_{l}[t]$[25]. This approach facilitates identity mapping, mitigates the vanishing/exploding gradient problem and achieves higher accuracy with fewer time steps compared to existing SNN models.
SEW ResNet [25] was also the first work to propose and apply a surrogate gradient approach for optimizing SNNs. Using backpropagation methods to train SNNs [29, 13] has become a common practice in neuromorphic computing. However, applying backpropagation in SNNs presents certain challenges. In SNNs, spike signals are typically modeled using binary values (0s and 1s) from a Heaviside function $H(\theta,t)$, where $\theta$ is the membrane potential and $t$ is a threshold:
$H(\theta,t)=\begin{cases}1&\text{if $\theta-t\geq 0$}\\ 0&\text{otherwise}\end{cases}$ (2)
The derivative $\partial H$ of this function is given by:
@@
$\frac{\partial H}{\partial\theta}=\begin{cases}+\infty&\text{if $\theta-t=0$}\\ 0&\text{otherwise}\end{cases}$ (3)
As shown, directly differentiating $H(\theta,t)$ results in unstable gradients. To address this, [25] proposed substituting the derivative of the Heaviside function $\partial H$ with a numerically stable gradient from a different activation function during backpropagation, such as the derivative $\partial\sigma$ of the Sigmoid function $\sigma(\theta,t)$:
$\sigma(\theta,t)=\frac{1}{1+e^{\theta-t}}$ (4)
$\frac{\partial H}{\partial\theta}=\frac{\partial\sigma}{\partial\theta}=\frac{e^{\theta-t}}{(e^{\theta-t}+1)^{2}}$
This approach yields a more stable gradient during backpropagation, allowing the use of existing automatic differentiation frameworks to optimize SNNs effectively.
With these advantages, we decided to choose SEW Resnet[25] architecture for our study.
III Bit Planes and Color Models in Deep Learning
@@
This section presents an overview of bit planes and color models, along with their applications in deep learning algorithms.
III-A Bit planes in deep learning
@@
Bit planes have shown extensive utility in deep learning applications, offering new ways to improve model efficiency and performance. In [10], researchers used bit planes to encode the input layer in Binary Neural Network (BNN). They argued that bit planes retain crucial spatial information, enabling a reduction in the number of multiply-accumulate operations (MACs) without compromising the model’s accuracy or robustness. This demonstrates the potential of bit planes in streamlining computations while maintaining competitive performance.
@@
The use of bit plane slicing on RGB images for adversarial robustness was explored in [30]. By selecting the most significant bit planes, the authors developed a classification model that was better equipped to handle adversarial attacks. This approach capitalizes on the spatial information encoded in different bit planes, which helps the model resist manipulations that would otherwise degrade its performance.
@@
Although bit planes have been widely studied and applied in ANNs, their utilization in SNNs is still unexplored. As far as we concern, our research is first to explore application of bit planes on SNNs spike coding process.
III-B Color models in deep learning
Color models are fundamental in computer graphics as they define how colors are represented, manipulated, and displayed for digital images. In computer graphics applications, understanding and effectively managing color models is crucial for tasks such as rendering realistic scenes, creating digital artwork, and designing user interfaces. Color models can be divided into device-oriented models, user-oriented models and device-independent models [32, 33].
Device-oriented color models are influenced by the signals of the devices and the resulting colors are affected by the technology used for display. Widely utilized in various applications that require color consistency with hardware tools, examples include any hardware devices used for human visual perception, such as televisions and video systems. Popular color models for devices include RGB, CMY(K) and YCbCr.
User-oriented color models are considered as a pathway between the observer and the device handling the color information. For our study, we mostly investigate cylindrically described models, including HSL and HSV. According to [32], these models also inherit other special properties such as device dependence and cylindrical-coordinate representation.
Device-independent color models are essential for defining color signals without reliance on the specific characteristics of any particular device or application. These models play a crucial role in applications that involve color comparisons and the transmission of visual information across networks that connect various hardware platforms. We will mostly research color models defined by the International Commission on Illumination (CIE), including CIE XYZ and CIE LAB.
Color models are becoming increasingly important in deep learning algorithms, offering valuable information that enhances model performance in various tasks. For example, in [34], a research on traffic light recognition tested a range of color models across different deep learning architectures, with results indicating that the RGB model remains favored despite some of its limitations. This suggests that, while RGB is not perfect, it still outperforms other models in certain applications.
In another study on applying convolutional neural networks (CNNs) for skin detection against adversarial attacks of generative models[35], RGB and YCbCr color models demonstrated superior performance, particularly in handling varying lighting conditions. This highlights the adaptability of these color models in scenarios where lighting inconsistencies can degrade model accuracy.
Moreover, the development of ColorNet[11] illustrates the potential of leveraging multiple color models to capture a broader set of features. By combining information from these diverse color models, ColorNet outperforms traditional models that rely on a single color model. However, this approach introduces computational overhead due to the inclusion of smaller submodels, which can be a challenge despite the model’s increased efficiency in some cases.
Despite extensive researches on the use of color models in ANN, their application in SNN remains underexplored and requires further investigation.
IV Proposed Method
In this section, we present in detail the design of our coding algorithm.
IV-A Hypothesis and Preliminary Study
@@
We hypothesize that the information contained in the bit planes of an image can be effectively used to encode inputs for SNN training. Bit planes decompose an image into binary levels, each representing a specific bit of the binary representation of pixel values. This decomposition can potentially preserve critical information of the original image while providing a format that might be more suitable for the spike-based processing in SNNs. The granularity of information available in bit planes offers a unique approach to encoding visual data, which could enhance the learning capabilities of SNNs by providing a richer, multi-level representation of the input.
To test this hypothesis, we conducted a preliminary experiment and analysis using the MNIST[36], which is a simple dataset containing gray-scale images of hand-written numbers. For analysis in term of information contained in the coded spike signals, we perform the calculation of the mean base-2 Shannon entropy[37] $\overline{H}(X_{j})$ as:
$\overline{H}(X_{j})=-\frac{1}{N}\sum_{j=1}^{N}\sum_{i}p(x_{i})\log_{2}p(x_{i})$ (5)
and the approximated average Kolmogorov complexity $\overline{K}$ across all dataset’s elements where the Kolmogorov complexity $K(X)$ of a coded image is estimated with DEFLATE compression algorithm[38] $C_{DEFLATE}(X)$ as:
$\overline{K}(X_{j})$ $=\frac{1}{N}\sum_{j=1}^{N}K(X_{j})$ (6)
$\approx\frac{1}{N}\sum_{j=1}^{N}C_{DEFLATE}(X_{j})$
$=\frac{1}{N}\sum_{j=1}^{N}Huffman(LZ77(X_{j}))$
where $Huffman(X)$ is the Huffman coding[39] and $LZ77(X)$ is the LZ77 compression[40] (or sometimes refered to as sliding window compression).
We compared three approaches: traditional spike coding, bit plane coding only, and the combination of both spike coding and bit plane coding. The detailed settings of this experiment can be found in Section V-A. The results are provided in Table I.
Coding method Validation Accuracy (%) Averaged Kolmogorov complexity approximation (bits) Averaged Shannon entropy (bits)
Spike coding 98.62 721.5 0.53
Bit plane coding only 97.87 772.8 0.56
Combination of spike coding and bit plane coding 98.93 744.4 0.55
TABLE I: Average validation accuracy and analysis metrics comparison on MNIST[36]
The results demonstrate that the approximated Kolmogorov complexity and Shannon entropy of bit-plane-coded data are higher, suggesting that the information content is more complex and less predictable. This may help explain why the accuracy of models using exclusively bit-plane-encoded inputs is lower compared to the baseline. On the other hand, a combination of rate coding and bit-plane coding resulted in an increase in performance. This outcome demonstrates that bit planes contain valuable information that SNN can learn from to improve their classification ability. This encourages us to carry out a further study and perform testing on other datasets.
IV-B Proposed Coding Method with Bit Planes
@@
Figure 2: Our proposed coding method (lower) compare with traditional spike coding method (upper).
@@
As we can see in Figure 2, the proposed coding scheme integrates a dual approach featuring both traditional spike coding and bit plane coding methodologies. Initially, the process commences with the conventional spike coding procedure applied to the image tensor. The bit plane coding algorithm (Algorithm 1) is employed to transform an input image into binary planes. The final encoded tensors are then generated through tensor concatenation across the first dimension, merging the outputs from spike coding and bit plane coding. These encoded tensors are thereby prepared for utilization in inference tasks and/or training processes.
@@
Algorithm 1 is designed to encode an image tensor batch $X$ into $n_{bit}$ bit planes. To determine $n_{bit}$, it is essential to know the maximum possible value $X_{max}$ of the tensor batch. Since different color models have different $X_{max}$, this value must be predefined according to the selected color model. Once $X_{max}$ is defined, $n_{bit}$ can be calculate as $\left\lceil\log_{2}(X_{max})\right\rceil$.
@@
To calculate the bit planes of an unnormalized integer input tensor batch $X$, we perform element-wise division of $X$ by 2. The remainder from this division constitutes a bit plane of $X$, while the quotient is saved for the extraction of subsequent bit planes. This process is repeated until we obtain $n_{bit}$ bit planes. The resulting bit planes are then concatenated to form $n_{bit}$ binary tensor batch of $X_{shape}$. This format closely resembles that of rate coding and ensures determinism in the coding results without having to sample more than $n_{bit}$ times.
Symbol Meaning
$X$ Input image tensors
$X_{max}$ Integer value indicate the maximum value of input image tensor
$n_{bit}$ Minimum number of bit planes required to encode input $X$ into bit planes
$log_{2}()$ Binary logarithm function
$X_{shape}$ Dimensions of X, usually include number of images being processed $B$, number of channels $C$ and resolution $H,W$
$concat()$ Tensor concatenation function, concatenating a list of tensors resulted to a tensor contain spikes of shape ($n_{bit}$, $X_{shape}$)
$L$ A list-like [41] data structure, used to temporarily store coding result
$\mathbb{Z}$ The set of all integers
TABLE II: Table of symbols used in Algorithm 1.
Algorithm 1 Bit plane coding algorithm
$X_{max}$ is defined, $X_{max}>0$ and $X\in\mathbb{Z}$
function BitplaneEncode($X,X_{max}$)
@@
     $n_{bit}\leftarrow\left\lceil\log_{2}(X_{max})\right\rceil$triangleright Get number of bit required to encode image from highest possible value $X$
     $L\leftarrow[\text{ }]$triangleright Empty list initialization
     for each $X_{shape}.C$ do
@@
              $L.insert(X\mod{2})$triangleright Element-wise modulo of tensor, then store the bit plane to $L$
              $X\leftarrow\left\lfloor X/2\right\rfloor$triangleright Element-wise division of input image tensor and floor the result
@@
     end for
     return $concat(L)$triangleright Concatenate all stored sub tensors in $L$ to a single tensor
@@
V Experiments
V-A Experimental settings
To investigate performance of SNN for images, we employed the popular tool SEW-ResNet18 with ADD element-wise function [25]. The model and experiment process were implemented using PyTorch [42] and SpikingJelly[43]. Color conversion algorithms are based on implementations from Kornia [44]. For the baseline method in the experiments, the spike rate coding is employed with a step size of 10. To facilitate optimization, we utilize surrogate gradient methods applied to all IF neurons using the arctan surrogate function $S(U)$ [25, 13] with the defined derivative $\partial S$ of:
$\frac{\partial S}{\partial U}=\frac{1}{\pi}\cdot\frac{1}{1+(\pi U)^{2}}$ (7)
The optimization process is conducted with the Adam algorithm [45], using a learning rate of $lr=1\times 10^{-3}$ and parameters $\beta=(0.9,0.999)$ across all evaluated datasets. Cross-Entropy Loss with Softmax function and mean reduction $L(x,y)$[42] across $N$ mini-batch dimension was also widely employed as the loss function: